OpenStudy (anonymous):

Help Pleaseeeee!!!!!!! I will give medals and become fan! If a sample of barium contains 2.3 x 10^23 atoms of barium, how many grams of barium are in the sample

OpenStudy (anonymous):

oh and there is options for answers give me a sec to give them

OpenStudy (wolfe8):

So you know the number of atoms is the Avogadro number times the number of moles. And you know you can get the mass of material from the number of moles and relative molecular/atomic mass. Right?

OpenStudy (anonymous):

No i didn't, i am okay at chemistry, the only part that gets me is the math. I just need help working through this problem and a couple others, but you dont have to help with the others and here are the options a: 0.38g b: 2.62g c: 52.4g d: 359.5g

OpenStudy (wolfe8):

As I said, since atomic number is Avogadro number times number of moles, you divide it by A number to get # of moles. And then you multiply it by the RMM/RAM to get mass again.

OpenStudy (anonymous):

ok thank you i will try but i don't know h0w good i will be

OpenStudy (anonymous):

whoever looks at this i still need help!

OpenStudy (wolfe8):

I think the last picture I gave you says it all. It has all the steps you need.

OpenStudy (anonymous):

I fell very stupid because I still dont understand

OpenStudy (abb0t):

Trust me, the math is not that hard. If you go on, the math gets far worse. But anyways, start with what you're given: \(\sf \color{red}{2.3 \times 10^{23} ~atoms}\). Make sure that you note your units!

OpenStudy (abb0t):

Now, use avogadro's number: \(\sf \color{orange}{6.022 \times 10^{23}} \frac{mol}{atom}\) to divide and cancel out the units of atom that you have.

OpenStudy (agent0smith):

You have the number of atoms... divide that by a mole (6.02x10^23) to see how many moles, that's your first step... knowing the number of moles. THEN multiply the number of moles, by the atomic weight of barium (look at barium on a periodic table)

OpenStudy (abb0t):

You will be left with a small number. Next, look at your periodic table OR visit this website: http://www.ptable.com/ and look the element Ba (barium). multiply by the given mass on the p-table. Since you know that molar mass on the periodic table is in units of: \(\sf \color{blue}{\frac{grams}{mol}}\)

OpenStudy (abb0t):

You're plan should be: \(\sf \color{green}{atoms \rightarrow mol~of~Ba \rightarrow grams~of~Ba}\)

OpenStudy (agent0smith):

^yep. Basically you always want to get moles first, and use that to get whatever else you need (litres of gas, mass in grams, etc)

OpenStudy (anonymous):

thanks guys my internet crashed sorry....